# A hot spots theorem for the mixed eigenvalue problem with small Dirichet region

### v1

@@ Abstract.
We prove that on convex domains, first mixed Laplace eigenfunctions have no interior critical points if the Dirichlet region is connected and sufficiently small. We use this result to construct a new family of polygonal domains for which Rauch’s hot spots conjecture holds and to prove a new general theorem regarding the hot spots conjecture.
@@ 1. Introduction
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a bounded, convex Lipschitz domain. Let $D\subseteq\overline{\Omega}$ (where $\overline{\Omega}$ is the closure of $\Omega$ in ${\mathbb{R}}^{2}$) be a compact set. Consider the eigenvalue problem
@@
This eigenvalue problem admits a discrete, increasing sequence of non-negative eigenvalues $\{\lambda_{n}^{D}\}$ and a corresponding orthonormal basis of eigenfunctions $\{u_{n}^{D}\}$ for $L^{2}(\Omega)$ that (weakly) solve (1). When we wish to emphasize the dependence on the domain, we may write $\lambda_{n}^{D}(\Omega):=\lambda_{n}^{D}$. Note that because we did not place strong restrictions on $D$, the (weak) eigenfunctions may not vanish everywhere on $D$ (see [FNO21]). We will suppose throughout the paper that $D$ is chosen such that eigenfunctions of (1) vanish everywhere on $D$. For example, this holds if $D$ is the closure of a relatively open subset of $\partial\Omega$ or if $D$ is the closure of a Lipschitz subdomain of $\Omega$.
For $\epsilon>0$, let $B(0,\epsilon)$ be the disk of radius $\epsilon$ centered at the origin in ${\mathbb{R}}^{2}$.
@@
There exists $\epsilon>0$ such that if $D\subseteq B(0,\epsilon)$ and $D$ is connected, then $u_{1}^{D}$ has no critical points in $\Omega\setminus D$ [^1] Recall that a domain is, by definition, open, so the theorem does not prohibit critical points in $\partial\Omega$.
@@
Our proof of Theorem 1.1 is heavily inspired by Miyamoto’s proof of Theorem A in [M09].
@@
In the past year or so, there has been an increased focus by researchers on variants of the hot spots conjecture for the mixed problem. In 2024 alone, the following pre-prints have appeared: [H24a], [LY24], [AR24], and [H24b]. To our knowledge these are the first papers on the topic since the early 2000s. See [AR24] for an overview of earlier results.
Theorem 1.1 is the combination of Propositions 1.5 and 2.1 below.
@@
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a simply connected, bounded Lipschitz domain. Let $D\subseteq\overline{\Omega}$ be as above. For each $y\in\partial\Omega$, let $\nu(y)$ be the outward normal vector to the boundary at $y$. We say that the pair $(\Omega,D)$ is Neumann convex if for any $x\in\Omega$ and $y\in\partial\Omega\setminus D$, the dot product $(y-x)\cdot\nu(y)$ is non-negative.
@@
Suppose that $(\Omega,D)$ is Neumann convex and that $D$ is connected. Let $d$ be the diameter of $\Omega$, and let $j_{0}$ be the first zero of the Bessel function $J_{0}(x)$. If
@@
In Section 4, we show some explicit examples where this inequality holds, and we use these results to prove two new results on the hot spots conjecture. Let $P_{n}$ denote a regular polygon with $n$ edges for $n\geq 3$. Let $e$ be the interior of an edge of $P_{n}$, and let $Q_{n}$ be the reflection of $P_{n}$ over $e$. Let $K_{n}=P_{n}\cup e\cup Q_{n}$. Then we have
@@
Let $\Omega$ be a convex Lipschitz domain, and let $u$ be a second Neumann eigenfunction of $\Omega$. Let $\Omega_{+}=\{x\in\Omega\mid u(x)>0\}$. If the diameter of $\Omega_{+}$ is at most half of the diameter of $\Omega$, then $u$ has no interior critical points in $\Omega_{+}$. The same statement applies to the set on which $u$ is negative.
@@ 2. First mixed eigenvalue limits
By translating the domain $\Omega$, we may assume that the origin is contained in $D$. As in Section 1, let $d$ denote the diameter of $\Omega$:
$d=\text{diam}(\Omega)=\sup\{|x-y|:x,y\in\Omega\}.$
Let $V$ be the area of $\Omega\setminus B(0,d/4)$. Since $d$ is the diameter of $\Omega$, we have $V>0$.
@@
Suppose that $D_{\epsilon}\subseteq B(0,\epsilon)$ for some $40<\epsilon<d/$ 4. Then we have
$\lambda_{1}^{D_{\epsilon}}\leq\frac{2\pi\ln(d/(4\epsilon))}{V\ln(4\epsilon/d)^{2}}.$
In particular, $\lambda_{1}^{D_{\epsilon}}$ tends to zero as $\epsilon$ tends to zero.
@@
For $40<\epsilon<d/$ 4, define a test function $\varphi$ in polar coordinates by
$\varphi_{\epsilon}(r)=\begin{cases}0\;\;&\text{if}\;\;r\leq\epsilon\\ 1-\frac{1}{\ln(4\epsilon/d)}\ln(4r/d)\;\;&\text{if}\;\;\epsilon<r<d/4\\ 1\;\;&\text{if}\;\;r\geq d/4.\end{cases}$
Note that $\varphi_{\epsilon}$ is an element of $H^{1}_{D_{\epsilon}}(\Omega)$ and is thus a valid test function for the variational formula for $\lambda_{1}^{D_{\epsilon}}$. We bound the Rayleigh quotient of $\varphi_{\epsilon}$ by bounding $\int_{\Omega}|\nabla\varphi_{\epsilon}|^{2}$ from above and $\int_{\Omega}\varphi_{\epsilon}^{2}$ from below. The quotient of these estimates will yield the desired estimate on $\lambda_{1}^{D_{\epsilon}}$. First,
@@
$=\frac{2\pi\ln(d/(4\epsilon))}{\ln(4\epsilon/d)^{2}}.$
For the other bound, we simply use that $V>0$:
$\int_{\Omega}\varphi_{\epsilon}^{2}\geq\int_{\Omega\setminus B(0,d/4)}\varphi_{\epsilon}^{2}=V.$
@@
Our proof of Proposition 2.1 actually works without most of the restrictions on $\Omega$ and $D$ given in Section 1. For example, our proof applies to the case of any bounded Lipschitz domain $\Omega$. The proof also does not require $D$ to be connected or to be chosen such that eigenfunctions of (1) vanish pointwise on $D$.
@@
The fact that $\lambda_{1}^{D_{\epsilon}}$ tends to 00 as $\epsilon$ tends to zero is not new. One can find this fact, for instance, in [FNO21]. However, the authors in [FNO21] prove a significantly more precise result and state it in terms of the Sobolev capacity of $D$, so we chose to include a simple proof in this paper.
@@ 3. Proof of the main theorem
Suppose toward a contradiction that $u_{1}^{D}$ has a critical point $p\in\Omega$. Since $u$ does not vanish anywhere in $\Omega\setminus D$, we may suppose without loss of generality that $u\geq 0$. By translating $\Omega$, we may suppose without loss of generality that $p$ equals the origin. Using polar coordinates, let $w(x,y)=J_{0}(\sqrt{\lambda_{1}^{D}}r)$. For $(x,y)\in\overline{\Omega}$, we have $r<d$ (with strict inequality since $p$ is an interior point), from which it follows that $w>0$ in $\overline{\Omega}$. By definition, $w$ satisfies $-\Delta w=\lambda_{1}^{D}w$ in $\Omega$. Define $f(x,y)=u_{1}^{D}(0)w(x,y)-u_{1}^{D}(x,y)$. This is a Laplace eigenfunction with eigenvalue $\lambda_{1}^{D}$. Moreover, $f$ is non-constant since $\lambda_{1}>0$ and $f>0$ on $D$. We will use $f$ to construct a test function that contradicts the variational formulation (2) for $\lambda_{1}^{D}$.
We have $f(0)=0$ and $\nabla f(0)=0$ since 00 is a critical point for both $u_{1}^{D}$ and $w$. Thus, there exists a neighborhood $U$ of 00 such that $U\setminus f^{-1}(\{0\})$ has at least four connected components, with $f$ taking alternating signs on each connected component. By Lemma 3.1 and using that $f>0$ on $D$, $\Omega\setminus f^{-1}(\{0\})$ has at least four connected components, and $f$ is positive in at least two of these components. Since $D$ is connected, there exists a unique connected component of $\Omega\setminus f^{-1}(\{0\})$ whose closure intersects $D$. Thus, there exists a connected component $B$ of $\Omega\setminus f^{-1}(\{0\})$ on which $f$ is positive and whose closure does not intersect $D$. Let $\varphi=f\chi_{B}\in H^{1}_{D}(\Omega)$.
@@
By Proposition 1.5, we need only show that for sufficiently small $\epsilon$ and appropriate $D_{\epsilon}\subseteq B(0,\epsilon)$,
$\lambda_{1}^{D_{\epsilon}}\leq\Big{(}\frac{j_{0}}{d}\Big{)}^{2}.$
This follows from Proposition 2.1. ∎

### v2

# A hot spots theorem for the mixed eigenvalue problem with small Dirichlet region

@@ Abstract.
We prove that on convex domains, first mixed Laplace eigenfunctions have no interior critical points if the Dirichlet region is connected and sufficiently small. We also find two seemingly new estimates on the first mixed eigenvalue to give explicit examples of when the Dirichlet region is sufficiently small.
@@ 1. Introduction
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a bounded Lipschitz domain. Let $D\subseteq\overline{\Omega}$ (where $\overline{\Omega}$ is the closure of $\Omega$ in ${\mathbb{R}}^{2}$) be a compact set such that $\Omega\setminus D$ is connected. Consider the eigenvalue problem
@@
When $D$ has non-empty interior relative to $\overline{\Omega}$, we interpret this problem as deleting the relative interior of $D$ from $\Omega$ and enforcing Dirichlet boundary conditions on $\partial D$. We will suppose that $D$ is chosen such that this eigenvalue problem admits a discrete, increasing sequence of non-negative eigenvalues $\{\lambda_{n}^{D}\}$ and a corresponding orthonormal basis of eigenfunctions $\{u_{n}^{D}\}$ for $L^{2}(\Omega\setminus\text{int}(D))$ that (weakly) solve (1). When we wish to emphasize the dependence on the domain, we may write $\lambda_{n}^{D}(\Omega):=\lambda_{n}^{D}$. Note that because we did not place strong restrictions on $D$, the (weak) eigenfunctions may not vanish pointwise on $D$ (see [FNO21]). We will assume throughout the paper that $D$ is chosen such that eigenfunctions of (1) vanish everywhere on $D$. For example, this holds if $D$ is the closure of a relatively open subset of $\partial\Omega$ or if $D\subseteq\Omega$ and $\Omega\setminus D$ is a Lipschitz domain. However, one can show that $D$ may be a somewhat more pathological set, such as a line segment in the interior of $\Omega$.
For some subset $E\subseteq{\mathbb{R}}^{2}$, define the diameter of $E$ to be the quantity
$\sup\{|x-y|:x,y\in E\}.$
@@
Suppose that $\Omega$ is convex. There exists $\epsilon>0$ such that if $D$ is connected with diameter at most $\epsilon$, then $u_{1}^{D}$ has no critical points in $\Omega\setminus D$ [^1] Recall that a domain is, by definition, open, so the theorem does not prohibit critical points in $\partial\Omega$.
@@
Our proofs of Theorems 1.1, 1.6, and 2.1 are heavily inspired by Miyamoto’s proof of Theorem A in [M09]. By extension, Theorems 1.6 and 2.1 are modifications of Weinberger’s proof of the maximality of the second Neumann eigenvalue of a Euclidean ball [W56].
@@
There has been recent interest by researchers on variants of the hot spots conjecture for the mixed problem. In 2024, the following pre-prints have appeared: [H24a], [LY24], [AR24], and [H24b]. To our knowledge, these are the first papers on the topic since a handful of results in the early 2000s appeared: [P02], [BP04], and [BPP04].
Theorem 1.1 is the combination of Propositions 1.5 and 2.2 below.
@@
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a simply connected, bounded Lipschitz domain. Let $D\subseteq\overline{\Omega}$ be as above. For each $y$ in the smooth part of $\partial\Omega$, let $\nu(y)$ be the outward normal vector to the boundary at $y$. We say that the pair $(\Omega,D)$ is Neumann convex if for any $x\in\Omega$ and for any $y\in\partial\Omega\setminus D$ for which $\nu(y)$ is defined, the dot product $(y-x)\cdot\nu(y)$ is non-negative. Equivalently, $(\Omega,D)$ is Neumann convex if and only if $\partial\Omega\setminus D$ is contained in the boundary of the convex hull of $\Omega$.
@@
Suppose that $(\Omega,D)$ is Neumann convex and that $D$ is connected. Let $d$ be the diameter of $\Omega$, and let $j_{0}\approx 2.4048$ be the first zero of the Bessel function $J_{0}(x)$. If
@@
In Sections 2 and 4, we solve two shape optimization questions for the mixed problem that appear to be new. Using these estimates, we show some explicit examples where this inequality holds, and we use these results to prove some new results on the hot spots conjecture. Roughly speaking, Theorem 1.6 applies when $D$ is long and narrow, while Theorem 2.2 applies when the diameter of $D$ is small.
Theorem 1.6 gives an explicit geometric condition on $(\Omega,D)$ such that the inequality in Proposition 1.5 holds. Let $(\Omega,D)$ be a Neumann convex pair embedded in the upper half plane $\{(x,y)\mid y\geq 0\}$ such that $D$ is contained in the closed second quadrant $\{(x,y)\mid x\leq 0,y\geq 0\}$. Let $\ell$ be the length of the projection of $\Omega$ onto the $y$-axis. Let $A$ be the area of the intersection of $\Omega$ with the first quadrant $\{(x,y)\mid x,y\geq 0\}$. Then we have
Theorem 1.6.
The first mixed eigenvalue satisfies
$\lambda_{1}^{D}\leq\Big{(}\frac{\pi\ell}{2A}\Big{)}^{2}.$
This inequality is sharp, and equality is uniquely acheived when $\Omega$ is the rectangle $(0,A/\ell)\times(0,\ell)$ with $D=\{0\}\times[0,\ell]$. Moreover, if
$\frac{d\ell}{A}\leq\frac{2j_{0}}{\pi}\approx 1.531,$
then $\lambda_{1}^{D}\leq(j_{0}/d)^{2}$, so first mixed eigenfunctions have no critical points in $\Omega\setminus D$.
Let $P_{n}$ denote a regular polygon with $n$ edges for $n\geq 3$. Let $e$ be the relative interior of an edge of $P_{n}$, and let $Q_{n}$ be the reflection of $P_{n}$ over $e$. Let $K_{n}=P_{n}\cup e\cup Q_{n}$. An application of Theorem 1.6 and Theorem 1.1 of [H24a] is the following:
@@
Let $\Omega$ be a convex domain, and let $u$ be a second Neumann eigenfunction of $\Omega$. Let $\Omega_{+}=\{x\in\Omega\mid u(x)>0\}$. If the diameter of $\Omega_{+}$ is at most half of the diameter of $\Omega$, then $u$ has no interior critical points in $\Omega_{+}$. The same statement applies to the set on which $u$ is negative.
Remark 1.9.
We do not know of any examples of a domain with a second Neumann eigenfunction satisfying the hypothesis of Theorem 1.8. It would be interesting to find out out whether or not such a pair exists.
@@ 2. First mixed eigenvalue limits
Let $B(0,R_{1})$ denote the ball in ${\mathbb{R}}^{2}$ centered at the origin with radius $R_{1}>0$. Consider the class of all pairs $(\Omega,D)$ as in Section 1 such that $D\subseteq B(0,R_{1})$ and such that the area of $\Omega\setminus B(0,R_{1})$ is some fixed positive number $V$. Theorem 2.1 below shows that the annulus of inner radius $R_{1}$ and area $V$ maximizes the first mixed eigenvalue over all such pairs. This estimate is particularly useful when $\Omega$ is a doubly connected domain with $D$ equal to its inner boundary.
The existing literature on eigenvalue optimization for the mixed problem is somewhat limited, but there do exist previous results in this area; see [N58], [PW61], [H62], and [B80]. Hersch’s result in [H62] solves a similar problem to Theorem 2.1 in which the area of $\Omega$ itself as well as the length of the inner boundary are both fixed quantities. Using simple estimates on the eigenvalues of annuli (see Equation (3) and Remark 2.3 below), Theorem 2.1 gives a smaller upper bound on the first mixed eigenvalue than Hersch’s theorem does when $R_{1}$ is small and $L$ is large.
For $0<R_{1}<R_{2}$, let $A(R_{1},R_{2})=\{R_{1}<|(x,y)|<R_{2}\}$ be the annulus with inner radius $R_{1}$ and outer radius $R_{2}$ centered at the origin. Let $\Omega$ be a bounded Lipschitz domain, and let $D\subseteq\overline{\Omega}$ be as in Section 1. Suppose that we may translate $\Omega$ and $D$ such that $D$ is contained in the closed ball $\overline{B(0,R_{1})}$. Let $\Omega_{+}=\Omega\setminus B(0,R_{1})$, and let
$R_{2}=\sqrt{R_{1}^{2}+|\Omega_{+}|/\pi},$
where $|\Omega_{+}|$ is the area of $\Omega_{+}$.
Theorem 2.1.
$\lambda_{1}^{D}(\Omega)\leq\lambda_{1}^{\partial B(0,R_{1})}(A(R_{1},R_{2}))$. Equality holds if and only if $\Omega=A(R_{1},R_{2})$ with $D=\partial B(0,R_{1})$.
Proof.
Let $\varphi$ be a non-negative first mixed eigenfunction for $A(R_{1},R_{2})$ with $D=\partial B(0,R_{1})$. Since the first mixed eigenvalue is simple and $\varphi$ does not change signs, $\varphi$ must be a radial function. We claim that $\varphi$ is strictly monotonic in the radial direction. Indeed, if not, then the set of critical points of $\varphi$ contains a circle of radius $R_{3}\in(R_{1},R_{2})$. The restriction of $\varphi$ to the annulus $A(R_{3},R_{2})$ is a non-constant, nowhere vanishing Neumann eigenfunction of $A(R_{3},R_{2})$. Such a function cannot exist, so the claim must hold.
Extend $\varphi$ to $\{|(x,y)|>R_{2}\}$ by setting it equal to its (constant) value on the outer boundary of $A(R_{1},R_{2})$. Extend $\varphi$ to vanish in the disk $\{|(x,y)|<R_{1}\}$. Note that $A(R_{1},R_{2})$ has area $|\Omega_{+}|$, so
$|A(R_{1},R_{2})\setminus\Omega_{+}|=|\Omega_{+}\setminus A(R_{1},R_{2})|.$
Then the estimates
$\int_{\Omega}|\nabla\varphi|^{2}=\int_{\Omega\cap A(R_{1},R_{2})}|\nabla\varphi|^{2}\leq\int_{A(R_{1},R_{2})}|\nabla\varphi|^{2}\;\;\text{and}$
$\int_{\Omega}\varphi^{2}=\int_{\Omega\cap A(R_{1},R_{2})}\varphi^{2}+\int_{\Omega\setminus A(R_{1},R_{2})}\varphi^{2}\geq\int_{\Omega\cap A(R_{1},R_{2})}\varphi^{2}+\int_{A(R_{1},R_{2})\setminus\Omega}\varphi^{2}=\int_{A(R_{1},R_{2})}\varphi^{2}$
combine to give
$\lambda_{1}^{D}(\Omega)\leq\frac{\int_{\Omega}|\nabla\varphi|^{2}}{\int_{\Omega}\varphi^{2}}\leq\frac{\int_{A(R_{1},R_{2})}|\nabla\varphi|^{2}}{\int_{A(R_{1},R_{2})}\varphi^{2}}=\lambda_{1}^{\partial B(0,R_{1})}(A(R_{1},R_{2})).$
If $(\Omega,D)\neq(A(R_{1},R_{2}),\partial B(0,R_{1}))$, then the above inequalities become strict. ∎
@@
Suppose that $R_{1}\leq R_{2}/10$. Then
$\lambda_{1}^{D}(\Omega)\leq\frac{4}{R_{2}^{2}\ln(R_{2}/R_{1})}.$
In particular, for fixed $\Omega$, $\lambda_{1}^{D}(\Omega)$ approaches 00 as the diameter of $D$ approaches 00.
@@
We begin by bounding $\lambda_{1}^{\partial B(0,\epsilon)}(A(\epsilon,1))$ from above when $\epsilon<0.1$. Let $\varphi(r,\theta)=1-\ln(r)/\ln(\epsilon)$. Then $\varphi$ vanishes on $\{r=\epsilon\}$, so its restriction to the annulus is a valid test function. We have
$\int_{A(\epsilon,1)}|\nabla\varphi|^{2}=\frac{-2\pi}{\ln\epsilon}.$
Since $\epsilon<0.1$, we have $r\varphi^{2}>2r-1\geq 0$ for $r\in(1/2,1)$. Hence
@@
We then have
$\lambda_{1}^{\partial B(0,\epsilon)}(A(\epsilon,1))\leq\frac{-4}{\ln\epsilon}.$
Scaling a domain by a positive value $C$ rescales its eigenvalues by $1/C^{2}$. Thus, the more general annulus $A(R_{1},R_{2})$ satisfies the inequality
(3) $\lambda_{1}^{\partial B(0,R_{1})}(A(R_{1},R_{2}))=\frac{1}{R_{2}^{2}}\lambda_{1}^{\partial B(0,R_{1}/R_{2})}(A(R_{1}/R_{2},1))\leq\frac{4}{R_{2}^{2}\ln(R_{2}/R_{1})}.$
Theorem 2.1 then gives the desired estimate. ∎
@@
Using a Poincaré-type inequality, one can show that the first mixed eigenvalue of the annulus is bounded below by
$\lambda_{1}^{B(0,R_{1})}(A(R_{1},R_{2}))\geq\frac{2}{R_{2}^{2}\ln(R_{2}/R_{1})}.$
Hence, the estimate on $\lambda_{1}^{B(0,R_{1})}(A(R_{1},R_{2}))$ obtained in the proof of Proposition 2.2 is sharp up to a constant multiple. This should be compared with the more general result in [FNO21], which is stated in terms of the Sobolev capacity of $D$.
@@
The estimate on the first mixed eigenvalue of an annulus shows that $A(R_{1},R_{2})$ satisfies the inequality $\lambda_{1}^{\partial B(0,R_{1})}(A(R_{1},R_{2}))\leq(j_{0}/2R_{2})^{2}$ as long as
$R_{1}/R_{2}\leq e^{-16/j_{0}^{2}}\approx 0.0628.$
However, this is not optimal. Numerical results indicate that the largest $R_{1}/R_{2}$ for which $\lambda_{1}^{\partial B(0,R_{1})}(R_{1},R_{2})<(j_{0}/2R_{2})^{2}$ is between $0.1$ and $0.15$.
@@ 3. Proof of the main theorem
Suppose toward a contradiction that $u_{1}^{D}$ has a critical point $p\in\Omega$. Since $u$ does not vanish anywhere in $\Omega\setminus D$, we may suppose without loss of generality that $u_{1}^{D}\geq 0$. By translating $\Omega$, we may suppose without loss of generality that $p$ equals the origin. Using polar coordinates, let $w(x,y)=J_{0}(\sqrt{\lambda_{1}^{D}}r)$. For $(x,y)\in\overline{\Omega}$, we have $r<d$ (with strict inequality since $p$ is an interior point), from which it follows that $w>0$ in $\overline{\Omega}$. By definition, $w$ satisfies $-\Delta w=\lambda_{1}^{D}w$ in $\Omega$. Define $f(x,y)=u_{1}^{D}(0)w(x,y)-u_{1}^{D}(x,y)$. This is a Laplace eigenfunction with eigenvalue $\lambda_{1}^{D}$. Moreover, $f$ is non-constant since $\lambda_{1}>0$ and $f>0$ on $D$. We will use $f$ to construct a test function that contradicts the variational formulation (2) for $\lambda_{1}^{D}$.
We have $f(0)=0$ and $\nabla f(0)=0$ since 00 is a critical point for both $u_{1}^{D}$ and $w$. Thus, there exists a neighborhood $U$ of 00 such that $U\setminus f^{-1}(\{0\})$ has at least four connected components, with $f$ taking alternating signs on each connected component. By Lemma 3.1 and using that $f>0$ on $D$, $\Omega\setminus f^{-1}(\{0\})$ has at least four connected components, and $f$ is positive in at least two of these components. Since $D$ is connected, there exists a unique connected component of $\Omega\setminus f^{-1}(\{0\})$ whose closure contains $D$. Thus, there exists a connected component $B$ of $\Omega\setminus f^{-1}(\{0\})$ on which $f$ is positive and whose closure does not intersect $D$. Let $\varphi=f\chi_{B}\in H^{1}_{D}(\Omega)$.
@@
Combine Propositions 1.5 and 2.2. ∎